# Gaussian Process Learning of Nonlinear Dynamics

### v1

@@ 2.3 Bayesian predictions
$p({\bf\it\theta}_{1},\cdots,{\bf\it\theta}_{N}|\mathbf{d}_{1},\mathbf{u}_{1},\cdots,\mathbf{d}_{N},\mathbf{u}_{N})=\prod_{i=1}^{N}p\left({\bf\it\theta}_{i}|~{}\mathbf{d}_{i}=f_{i}(\mathbf{U};{\bf\it\theta}_{i}),\mathbf{u}_{i}\right)$ (14)
@@
[…]
in which all ${\bf\it\theta}_{i}$’s are marginalized. For instance, Monte Carlo sampling can be employed over the posterior $p({\bf\it\theta}_{1},\cdots,{\bf\it\theta}_{N}|\mathbf{d}_{1},\mathbf{u}_{1},\cdots,\mathbf{d}_{N},\mathbf{u}_{N})$ to estimate the mean function and second-order moments of ${\bf\it x}(t)$. In fact, $p({\bf\it x}(t)|{\bf\it\theta}_{1},\cdots,{\bf\it\theta}_{N})$ represents a deterministic dynamical model, and one can hence write the state solution ${\bf\it x}(t)$ determined by the parameters $({\bf\it\theta}_{1},\cdots,{\bf\it\theta}_{N})$ as a function of them, i.e., ${\bf\it x}(t;{\bf\it\theta}_{1},\cdots,{\bf\it\theta}_{N})$.
@@ 2.4 Implementation considerations
To further reduce the impact of data noise, we can employ a corrected version of $f_{i}(\hat{\mathbf{U}};{\bf\it\theta}_{i})$ with smoothed state data $\hat{\mathbf{U}}=[\hat{\mathbf{u}}_{i}~{}\cdots~{}\hat{\mathbf{u}}_{N}]^{\mathsf{T}}$ through Gaussian process reconstruction, i.e.,
@@
    Kernel parameters and noise term $\chi_{i}^{u}$: we simply consider the marginal likelihood function $p(\mathbf{u}_{i})$ for $\mathbf{u}_{i}$ only (with the $\mathbf{d}_{i}$ term marginalized). Given in the closed form
@@
    this marginal likelihood function only involves the kernel parameters and $\chi_{i}^{u}$ and can be maximized to determine their values.
@@
    Noise term $\chi_{i}^{d}$: the variance value $\chi_{i}^{d}$ for the additive noise onto the time derivatives is adjusted to ensure that $(\mathbf{R}_{i}^{dd})^{-1}=\partial_{t}\partial_{t^{\prime}}\kappa_{i}(\mathcal{T},\mathcal{T})+\chi^{d}_{i}\mathbf{I}_{K}-\mathbf{K}^{du}_{i}(\mathbf{K}^{uu}_{i})^{-1}\mathbf{K}^{ud}_{i}$ is well-conditioned, which is a numerical requirement for evaluating $\mathbf{R}_{i}^{dd}$ needed in the Bayesian inference. On the other hand, as $\chi_{i}^{d}$ measures the discrepancy between $\mathbf{d}_{i}$ and $f_{i}(\mathbf{U};{\bf\it\theta}_{i})$, it is a valid assumption that $\chi_{i}^{d}$ should not be a large value when the smoothed $f_{i}(\hat{\mathbf{U}};{\bf\it\theta}_{i})$ is used, for which we suggest $\chi_{i}^{d}$ and $\chi_{i}^{u}$ share the same order of magnitude.
@@ 3.1 Scenario (I): linear parametrization
in which ${\bf\it g}_{i}:\mathbb{R}^{N}\to\mathbb{R}^{p_{i}}$ is a given vector-valued function. $f_{i}(\mathbf{U};{\bf\it\theta}_{i})$ is hence rewritten as $\mathbf{G}_{i}{\bf\it\theta}_{i}$ with $\mathbf{G}_{i}:=[{\bf\it g}_{i}({\bf\it\hat{x}}(t_{0}))~{}\cdots~{}{\bf\it g}_{i}({\bf\it\hat{x}}(t_{K-1}))]^{\mathsf{T}}\in\mathbb{R}^{K\times p_{i}}$, a linear transformation acting on the parameter vector ${\bf\it\theta}_{i}$. Thus, the posterior (10) given by the proposed inference method is explicitly represented as
$\begin{split}p({\bf\it\theta}_{i}|~{}\mathbf{d}_{i}=\mathbf{G}_{i}{\bf\it\theta}_{i},\mathbf{u}_{i})&\propto\exp\left(-\frac{1}{2}\left(\left\|\mathbf{G}_{i}{\bf\it\theta}_{i}-\hat{\mathbf{d}}_{i}\right\|_{\mathbf{R}_{i}^{dd}}^{2}+\|{\bf\it\theta}_{i}\|_{{\bf\it\Lambda}_{i}}^{2}\right)\right)\\ &\propto\exp\left(-\frac{1}{2}({\bf\it\theta}_{i}-{\bf\it\mu}_{i})^{\mathsf{T}}{\bf\it\Sigma}_{i}^{-1}({\bf\it\theta}_{i}-{\bf\it\mu}_{i})\right)\,,\end{split}$ (19)
@@
In the Bayesian inference of 19, each penalty coefficient $\lambda_{ij}$ in ${\bf\it\Lambda}_{i}$ is inversely proportional to the variance of the corresponding Gaussian prior for $\theta_{ij}$ ($j$th entry of ${\bf\it\theta}_{i}$). In other words, the magnitude of $\lambda_{ij}$ reflects the ‘confidence’ in the given prior, and this will further influence the Bayesian estimation, especially the reflection of sparsity promotion in uncertainty quantification. In this paper, we suggest the following steps to set $\lambda_{ij}$ values for each equation $i$:
@@ 4.1 Lotka–Volterra equations
For data generation, the dynamical system is solved using the forward Euler method with time step $\Delta t=0.001$ over $[0,T]$ where $T=20$, and the initial condition is $x_{1}(0)=x_{2}(0)=1$. The generated data are split into two phases, $[0,0.4T]$ for training and $(0.4T,T]$ for testing/prediction. While the step size $\Delta t$ is chosen to ensure the stability and accuracy of time integration, such data density is far beyond needed for model training. We consider that only $25\%$ randomly sampled data from those collected over $[0,0.4T]$ are used for training. This subset is treated as the entire dataset available, i.e., the case of $100\%$ data in the upcoming results. To investigate the impact of data sparsity on parameter estimation and uncertainty quantification, we further reduce the data density to $10\%$, $5\%$, and even $1\%$. Meanwhile, we introduce three levels of white noise – $0\%,10\%$, and $20\%$, respectively – to the training data. The noise term follows a normal distribution with zero mean, and its standard deviation is defined by multiplying the average of data by the noise level.
We are going to test the two cases in Scenario (I) introduced in Subsection 3.1.
@@
noise free parameters 100% data 10% data 5% data 1% data ground truth $\alpha$ 1.500 ±$\pm~{}(2.19\times 10^{-4})$ 1.500 ±$\pm~{}(7.46\times 10^{-4})$ 1.500 ±$\pm~{}(1.29\times 10^{-3})$ 1.524 ±$\pm~{}(8.04\times 10^{-3})$ 1.5 $\beta$ 1.000 ±$\pm~{}(1.34\times 10^{-4})$ 1.000 ±$\pm~{}(4.63\times 10^{-4})$ 0.998 ±$\pm~{}(9.29\times 10^{-4})$ 1.050 ±$\pm~{}(1.41\times 10^{-2})$ 1 $\delta$ 1.000 ±$\pm~{}(2.15\times 10^{-4})$ 1.000 ±$\pm~{}(7.40\times 10^{-4})$ 0.998 ±$\pm~{}(1.30\times 10^{-3})$ 0.955 ±$\pm~{}(8.12\times 10^{-3})$ 1 $\gamma$ 3.000 ±$\pm~{}(1.33\times 10^{-4})$ 3.000 ±$\pm~{}(4.64\times 10^{-4})$ 2.996 ±$\pm~{}(9.28\times 10^{-4})$ 2.934 ±$\pm~{}(1.43\times 10^{-2})$ 3 $10\%$ noise level parameters 100% data 10% data 5% data 1% data ground truth $\alpha$ 1.507 ±$\pm~{}(6.55\times 10^{-3})$ 1.484 ±$\pm~{}(2.21\times 10^{-2})$ 1.448 ±$\pm~{}(3.17\times 10^{-2})$ 1.583 ±$\pm~{}(1.24\times 10^{-1})$ 1.5 $\beta$ 1.002 ±$\pm~{}(4.03\times 10^{-3})$ 0.977 ±$\pm~{}(1.35\times 10^{-2})$ 0.934 ±$\pm~{}(1.95\times 10^{-2})$ 1.254 ±$\pm~{}(1.20\times 10^{-1})$ 1 $\delta$ 1.001 ±$\pm~{}(6.49\times 10^{-3})$ 0.996 ±$\pm~{}(2.19\times 10^{-2})$ 0.990 ±$\pm~{}(3.19\times 10^{-2})$ 0.804 ±$\pm~{}(1.25\times 10^{-1})$ 1 $\gamma$ 2.989 ±$\pm~{}(3.99\times 10^{-3})$ 2.991 ±$\pm~{}(1.35\times 10^{-2})$ 2.969 ±$\pm~{}(1.96\times 10^{-2})$ 2.344 ±$\pm~{}(1.21\times 10^{-1})$ 3 $20\%$ noise level parameters 100% data 10% data 5% data 1% data ground truth $\alpha$ 1.509 ±$\pm~{}(1.30\times 10^{-2})$ 1.472 ±$\pm~{}(4.48\times 10^{-2})$ 1.389 ±$\pm~{}(6.30\times 10^{-2})$0.050 ±$\pm~{}(2.25\times 10^{-1})$ 1.5 $\beta$ 1.000 ±$\pm~{}(8.01\times 10^{-3})$ 0.954 ±$\pm~{}(2.68\times 10^{-2})$ 0.878 ±$\pm~{}(3.80\times 10^{-2})$ 0.019 ±$\pm~{}(1.83\times 10^{-1})$ 1 $\delta$ 1.003 ±$\pm~{}(1.29\times 10^{-2})$ 0.977 ±$\pm~{}(4.43\times 10^{-2})$ 0.987 ±$\pm~{}(6.20\times 10^{-2})$ 0.299 ±$\pm~{}(2.21\times 10^{-1})$ 1 $\gamma$ 2.979 ±$\pm~{}(7.96\times 10^{-2})$ 2.934 ±$\pm~{}(2.70\times 10^{-2})$ 2.976 ±$\pm~{}(3.73\times 10^{-2})$ 1.025 ±$\pm~{}(1.83\times 10^{-1})$ 3
@@
Table 1: Mean $\pm$ standard deviation of estimated model parameters with different noise levels and data densities.
@@
Figure 1: Trajectory reconstruction ($t\in[0,8]$) and prediction ($t\in(8,20]$) of prey and predator dynamics with estimated parameters at $0\%$, $10\%$, and $20\%$ noise levels using $100\%$, $10\%$, $5\%$ and $1\%$ data. The uncertainty bands denote ±$\pm~{}\sigma$ (standard deviation) of the Bayesian prediction. The grey vertical line separates the training and testing stages.
@@
The posterior mean and standard deviation of ${\bf\it\theta}_{1}$ and ${\bf\it\theta}_{2}$ estimated under various conditions are presented in Table 1. In the noise-free cases, the inferred parameters coincide with the ground truth, and the small standard deviations indicate low uncertainties in the estimation. With less training data, a minor reduction in the accuracy of the mean predictions is observed, accompanied by naturally expanded deviations. Nevertheless, even when utilizing only $1\%$ of the data, the inference result still aligns closely with the ground truth, and the corresponding uncertainty remains at the $10^{-2}$ level. In the noise-corrupted cases, the compromise of accuracy in parameter inference becomes more obvious, especially when the data density is diminished. Although the posterior mean values remain accurate with $100\%$ data, the standard deviations are increased by the corruption of data noise. The most challenging case arises when $1\%$ of the data with $20\%$ noise level are adopted. It leads to a complete breakdown of the parameter estimation.
The obtained posterior distributions of system parameters are subsequently employed in Bayesian prediction using 15. The predictions are visualized in Figure 1. The time window before the vertical gray line ($t=0.4T=8$) corresponds to the training data coverage, while the window afterwards is the prediction phase. Evidently, the existence of data noise significantly impacts the predictive uncertainty and requires a sufficient amount of data to guarantee accuracy. Solutions under two new initial conditions, respectively being $(x_{1}(0),x_{2}(0))=(2,10)$ and $(x_{1}(0),x_{2}(0))=(2,1.2)$, are evaluated numerically and shown in Figure 2. These two examples serve to demonstrate the fact that the learned dynamical model is (and should be) independent of the initial conditions of training trajectories. These solutions are respectively derived with $100\%$ and $10\%$ training data densities, both corrupted by $10\%$-level noise. The results present a close alignment between the predictions and ground truth of the trajectories.
Figure 2: Trajectory predictions for two new initial conditions issued by dynamical models respectively trained with $100\%$ and $10\%$ of data, both corrupted by $10\%$ noise.
@@
Figure 3: Posterior distributions of ${\bf\it\theta}_{i}$ estimated under 0% and 20% noise levels with 10% and 50% data.
@@
The relative error in the posterior means and their relative deviation of the estimated parameters at different noise levels and amounts of training data, denoted by $\epsilon_{1}$ and $\epsilon_{2}$ respectively, are defined as follows and shown in Figure 4:
$\epsilon_{1}=100\%\times\left(\frac{\sum_{i=1}^{2}\|{\bf\it\theta}^{*}_{i}-\mathbb{E}_{\text{post}}[{\bf\it\theta}_{i}]\|_{2}^{2}}{\sum_{i=1}^{2}\|{\bf\it\theta}^{*}_{i}\|_{2}^{2}}\right)^{1/2}\,,\quad\text{and}\quad\epsilon_{2}=100\%\times\left(\frac{\sum_{i=1}^{2}\mathbb{V}ar_{\text{post}}[{\bf\it\theta}_{i}]}{\sum_{i=1}^{2}\|{\bf\it\theta}^{*}_{i}\|_{2}^{2}}\right)^{1/2}\,,$ (25)
in which $\mathbb{E}_{\text{post}}[{\bf\it\theta}_{i}]$ and $\mathbb{V}ar_{\text{post}}[{\bf\it\theta}_{i}]$ respectively represent the posterior means and variances of parameters ${\bf\it\theta}_{i}$, and ${\bf\it\theta}^{*}_{i}$ is the ground truth (reference solution). It is evident that, overall, the values of both these error/uncertainty indicators are effectively decreased as the data amount is increased or the noise is reduced, i.e., an convergence is confirmed by these observations.
@@
Results with scarce, noise-free data, including the trajectory reconstruction and prediction, as well as the approximation of $f(\cdot;{\bf\it\theta})$, are presented in Figure 5. The neural network approximates $f(\cdot;{\bf\it\theta})$ well within the range covered by training data, but fails in extrapolation when the prediction of $x$ goes beyond the training coverage. In fact, this observation is aligned with the limitation of neural network regression in generalization. Unlike the previously discussed scenarios of parameter estimation and sparse identification, the neural network approximation may not recover the actual underlying dynamical system but only provides a reparametrization. Therefore, the poor predictive performance beyond training coverage is reasonable and expected. Nevertheless, such inadequacy in predictions can be effectively indicated by the uncertainty bands in our results. On the other hand, if the training data have covered the entire time domain of interest, the neural network reparametrization $f(\cdot;{\bf\it\theta})$ serves a surrogate model. As shown in Figure 6, only with 10 data points, the surrogate modeling is already decent and enables a good trajectory reconstruction.

### v2

@@ 2.3 Bayesian predictions
$\pi_{\texttt{post}}({\bf\it\theta}_{1},\cdots,{\bf\it\theta}_{N})=p\left({\bf\it\theta}_{1},\cdots,{\bf\it\theta}_{N}\mid\left\{{\bf\it\delta}_{i}=\mathbf{0},~{}{\bf\it U}_{i}=\mathbf{u}_{i}\right\}_{i=1}^{N},\mathbf{Z}\right)=\prod_{i=1}^{N}\pi_{\texttt{post}}({\bf\it\theta}_{i})\,.$ (15)
@@
$p\left({\bf\it x}(t)\mid\left\{{\bf\it\delta}_{i}=\mathbf{0},~{}{\bf\it U}_{i}=\mathbf{u}_{i}\right\}_{i=1}^{N},\mathbf{Z}\right)=\int p({\bf\it x}(t)|{\bf\it\theta}_{1},\cdots,{\bf\it\theta}_{N})\prod_{i=1}^{N}\pi_{\texttt{post}}({\bf\it\theta}_{i})\,,$ (16)
in which all ${\bf\it\theta}_{i}$’s are marginalized. For instance, Monte Carlo sampling can be employed over the posterior $\pi_{\texttt{post}}({\bf\it\theta}_{1},\cdots,{\bf\it\theta}_{N})$ to estimate the mean function and second-order moments of ${\bf\it x}(t)$. In fact, $p({\bf\it x}(t)|{\bf\it\theta}_{1},\cdots,{\bf\it\theta}_{N})$ represents a deterministic dynamical model, and one can hence write the state solution ${\bf\it x}(t)$ determined by the parameters $({\bf\it\theta}_{1},\cdots,{\bf\it\theta}_{N})$ as a function of them, i.e., ${\bf\it x}(t;{\bf\it\theta}_{1},\cdots,{\bf\it\theta}_{N})$.
@@ 2.4 Implementation considerations
To further reduce the impact of data noise, we can employ a corrected version of $f_{i}(\hat{\mathbf{Z}};{\bf\it\theta}_{i})$ with smoothed state data $\hat{\mathbf{Z}}=[\hat{\mathbf{u}}_{i}~{}\cdots~{}\hat{\mathbf{u}}_{N}]^{\mathsf{T}}$ through Gaussian process reconstruction, i.e.,
@@
    Kernel parameters and noise term $\chi_{i}^{u}$: we simply consider the marginal likelihood function $p(\mathbf{u}_{i})$ for ${\bf\it U}_{i}$ only (with the ${\bf\it D}_{i}$ term marginalized). Given in the closed form
@@
    this marginal likelihood function only involves the kernel parameters and $\chi_{i}^{u}$ and can be maximized to determine their values. It is worth pointing out that the Gaussian process hyperparameters are determined individually to adapt to the amplitude and lengthscale of each $x_{i}$, and thus the covariance is not constant across the components of the state vector ${\bf\it x}$.
@@
    Noise term $\chi_{i}^{d}$: the variance value $\chi_{i}^{d}$ for the additive noise onto the time derivatives is adjusted to ensure that $(\mathbf{R}_{i}^{dd})^{-1}=\partial_{t}\partial_{t^{\prime}}\kappa_{i}(\mathcal{T},\mathcal{T})+\chi^{d}_{i}\mathbf{I}_{K}-\mathbf{K}^{du}_{i}(\mathbf{K}^{uu}_{i})^{-1}\mathbf{K}^{ud}_{i}$ is well-conditioned, which is a numerical requirement for evaluating $\mathbf{R}_{i}^{dd}$ needed in the Bayesian inference. On the other hand, as $\chi_{i}^{d}$ measures the discrepancy between $\mathbf{d}_{i}$ and $f_{i}(\mathbf{Z};{\bf\it\theta}_{i})$, it is a valid assumption that $\chi_{i}^{d}$ should not be a large value when the smoothed $f_{i}(\hat{\mathbf{Z}};{\bf\it\theta}_{i})$ is used, for which we suggest $\chi_{i}^{d}$ and $\chi_{i}^{u}$ share the same order of magnitude.
@@ 3.1 Scenario (I): linear parametrization
in which ${\bf\it g}_{i}:\mathbb{R}^{N}\to\mathbb{R}^{p_{i}}$ is a given vector-valued function. $f_{i}(\mathbf{Z};{\bf\it\theta}_{i})$ is hence rewritten as $\mathbf{G}_{i}{\bf\it\theta}_{i}$ with $\mathbf{G}_{i}:=[{\bf\it g}_{i}({\bf\it\hat{x}}(t_{0}))~{}\cdots~{}{\bf\it g}_{i}({\bf\it\hat{x}}(t_{K-1}))]^{\mathsf{T}}\in\mathbb{R}^{K\times p_{i}}$, a linear transformation acting on the parameter vector ${\bf\it\theta}_{i}$. Thus, the posterior (11) given by the proposed inference method is explicitly represented as
$\begin{split}\pi_{\texttt{post}}({\bf\it\theta}_{i})&\propto\exp\left(-\frac{1}{2}\left(\left\|\mathbf{G}_{i}{\bf\it\theta}_{i}-\hat{\mathbf{d}}_{i}\right\|_{\mathbf{R}_{i}^{dd}}^{2}+\|{\bf\it\theta}_{i}\|_{{\bf\it\Lambda}_{i}}^{2}\right)\right)\\ &\propto\exp\left(-\frac{1}{2}({\bf\it\theta}_{i}-{\bf\it\mu}_{i})^{\mathsf{T}}{\bf\it\Sigma}_{i}^{-1}({\bf\it\theta}_{i}-{\bf\it\mu}_{i})\right)\,,\end{split}$ (20)
@@
In the Bayesian inference of 20, each penalty coefficient $\lambda_{ij}$ in ${\bf\it\Lambda}_{i}$ is inversely proportional to the variance of the corresponding Gaussian prior for $\theta_{ij}$ ($j$th entry of ${\bf\it\theta}_{i}$). In other words, the magnitude of $\lambda_{ij}$ reflects the ‘confidence’ in the given prior, and this will further influence the Bayesian estimation, especially the reflection of sparsity promotion in uncertainty quantification. In this paper, we suggest the following steps to set $\lambda_{ij}$ values for each equation $i$:
@@ 4.1 Lotka–Volterra equations
For data generation, the dynamical system is solved using the implicit Euler method with time step $\Delta t=0.001$ over $[0,T]$ where $T=20$, and the initial condition is $x_{1}(0)=x_{2}(0)=1$. The generated data are split into two phases, $[0,0.4T]$ for training and $(0.4T,T]$ for testing/prediction. While the step size $\Delta t$ is chosen to ensure the stability and accuracy of time integration, such data density is far beyond needed for model training. We consider that only $25\%$ randomly sampled data from those collected over $[0,0.4T]$ are used for training. This subset is treated as the entire dataset available, i.e., the case of $100\%$ data (2000 points) in the upcoming results. To investigate the impact of data sparsity on parameter estimation and uncertainty quantification, we further reduce the data density to $10\%$, $5\%$, and even $1\%$, (i.e., 200, 100 and 20 data points, respectively). Meanwhile, we introduce three levels of white noise – $0\%,10\%$, and $20\%$, respectively – to the training data. The noise term follows a normal distribution with zero mean, and its standard deviation is defined by multiplying the average of data by the noise level.
We are going to test the two cases in Scenario (I) introduced in Subsection 3.1. When analyzing the results of parameter estimation, the overall relative error in the posterior means of the estimated parameters and their relative deviation, respectively denoted by $\epsilon_{1}$ and $\epsilon_{2}$, are defined as follows:
$\epsilon_{1}=100\%\times\left(\frac{\sum_{i=1}^{2}\|{\bf\it\theta}^{*}_{i}-\mathbb{E}_{\text{post}}[{\bf\it\theta}_{i}]\|_{2}^{2}}{\sum_{i=1}^{2}\|{\bf\it\theta}^{*}_{i}\|_{2}^{2}}\right)^{1/2}\,,\,\text{and}\,\epsilon_{2}=100\%\times\left(\frac{\sum_{i=1}^{2}\mathbb{V}ar_{\text{post}}[{\bf\it\theta}_{i}]}{\sum_{i=1}^{2}\|{\bf\it\theta}^{*}_{i}\|_{2}^{2}}\right)^{1/2}\,,$ (26)
in which $\mathbb{E}_{\text{post}}[{\bf\it\theta}_{i}]$ and $\mathbb{V}ar_{\text{post}}[{\bf\it\theta}_{i}]$ respectively represent the posterior means and variances of parameters ${\bf\it\theta}_{i}$, and ${\bf\it\theta}^{*}_{i}$ is the ground truth (reference solution).
@@
noise free parameters 100% data 10% data 5% data 1% data ground truth $\alpha$ 1.500 ±$\pm~{}(6.94\times 10^{-6})$ 1.500 ±$\pm~{}(2.36\times 10^{-4})$ 1.499 ±$\pm~{}(4.19\times 10^{-4})$ 1.511 ±$\pm~{}(2.81\times 10^{-3})$ 1.5 $\beta$ 1.001 ±$\pm~{}(4.29\times 10^{-5})$ 1.001 ±$\pm~{}(1.47\times 10^{-4})$ 1.000 ±$\pm~{}(3.03\times 10^{-4})$ 1.022 ±$\pm~{}(5.83\times 10^{-3})$ 1 $\delta$ 1.001 ±$\pm~{}(6.11\times 10^{-5})$ 1.001 ±$\pm~{}(2.10\times 10^{-4})$ 1.000 ±$\pm~{}(3.74\times 10^{-4})$ 0.975 ±$\pm~{}(5.27\times 10^{-3})$ 1 $\gamma$ 3.002 ±$\pm~{}(1.90\times 10^{-4})$ 3.003 ±$\pm~{}(6.41\times 10^{-4})$ 2.999 ±$\pm~{}(1.16\times 10^{-3})$ 2.964 ±$\pm~{}(8.42\times 10^{-3})$ 3 $10\%$ noise level parameters 100% data 10% data 5% data 1% data ground truth $\alpha$ 1.506 ±$\pm~{}(6.51\times 10^{-3})$ 1.486 ±$\pm~{}(2.21\times 10^{-2})$ 1.449 ±$\pm~{}(3.20\times 10^{-2})$ 1.579 ±$\pm~{}(1.25\times 10^{-1})$ 1.5 $\beta$ 1.003 ±$\pm~{}(4.03\times 10^{-3})$ 0.979 ±$\pm~{}(1.37\times 10^{-2})$ 0.937 ±$\pm~{}(1.97\times 10^{-2})$ 1.226 ±$\pm~{}(1.16\times 10^{-1})$ 1 $\delta$ 1.002 ±$\pm~{}(2.82\times 10^{-3})$ 0.996 ±$\pm~{}(8.50\times 10^{-3})$ 0.990 ±$\pm~{}(1.35\times 10^{-2})$ 0.792 ±$\pm~{}(6.45\times 10^{-2})$ 1 $\gamma$ 2.991 ±$\pm~{}(8.74\times 10^{-3})$ 2.992 ±$\pm~{}(2.66\times 10^{-2})$ 2.967 ±$\pm~{}(4.25\times 10^{-2})$ 2.322 ±$\pm~{}(1.65\times 10^{-1})$ 3 $20\%$ noise level parameters 100% data 10% data 5% data 1% data ground truth $\alpha$ 1.509 ±$\pm~{}(1.31\times 10^{-2})$ 1.474 ±$\pm~{}(4.45\times 10^{-2})$ 1.389 ±$\pm~{}(6.28\times 10^{-2})$ 1.593 ±$\pm~{}(2.13\times 10^{-1})$ 1.5 $\beta$ 1.001 ±$\pm~{}(8.06\times 10^{-3})$ 0.956 ±$\pm~{}(2.70\times 10^{-2})$ 0.881 ±$\pm~{}(3.79\times 10^{-2})$ 1.304 ±$\pm~{}(1.91\times 10^{-1})$ 1 $\delta$ 1.004 ±$\pm~{}(5.65\times 10^{-3})$ 0.976 ±$\pm~{}(1.72\times 10^{-2})$ 0.986 ±$\pm~{}(2.74\times 10^{-2})$ 0.758 ±$\pm~{}(1.17\times 10^{-1})$ 1 $\gamma$ 2.981 ±$\pm~{}(1.74\times 10^{-2})$ 2.933 ±$\pm~{}(5.38\times 10^{-2})$ 2.970 ±$\pm~{}(8.57\times 10^{-2})$ 2.050 ±$\pm~{}(2.99\times 10^{-1})$ 3
@@
Table 1: Mean $\pm$ standard deviation of estimated model parameters under 0%, 10% and 20% noise level with 100%, 10%, 5%, and 1% data densities (2000, 200, 100, and 20 data points, respectively)
@@
Figure 1: Trajectory reconstruction ($t\in[0,8]$) and prediction ($t\in(8,20]$) of prey and predator dynamics with estimated parameters at $0\%$, $10\%$, and $20\%$ noise levels using $100\%$, $10\%$, $5\%$, and $1\%$ data (2000, 200, 100, and 20 data points, respectively). The uncertainty bands denote ±$\pm~{}\sigma$ (standard deviation) of the Bayesian prediction. The grey vertical line separates the training and testing stages.
@@
The posterior mean and standard deviation of ${\bf\it\theta}_{1}$ and ${\bf\it\theta}_{2}$ estimated under various conditions are presented in Table 1. In the noise-free cases, the inferred parameters coincide with the ground truth, and the small standard deviations indicate low uncertainties in the estimation. With less training data, a minor reduction in the accuracy of the mean predictions is observed, accompanied by naturally expanded deviations. Nevertheless, even when utilizing only $1\%$ of the data, the inference result still aligns closely with the ground truth, and the corresponding uncertainty remains at the $10^{-3}$ level. In the noise-corrupted cases, the compromise of accuracy in parameter inference becomes more obvious, especially when the data density is diminished. Although the posterior mean values remain accurate with $100\%$ data, the standard deviations are increased by the corruption of data noise. The most challenging case arises when $1\%$ of the data with $20\%$ noise level are adopted. It leads to a complete breakdown of the parameter estimation.
The obtained posterior distributions of system parameters are subsequently employed in Bayesian prediction using 16. The predictions are visualized in Figure 1. The time window before the vertical gray line ($t=0.4T=8$) corresponds to the training data coverage, while the window afterwards is the prediction phase. Evidently, the existence of data noise significantly impacts the predictive uncertainty and requires a sufficient amount of data to guarantee accuracy. Solutions under two new initial conditions, respectively being $(x_{1}(0),x_{2}(0))=(2,10)$ and $(x_{1}(0),x_{2}(0))=(2,1.2)$, are evaluated numerically and shown in Figure 2. These two examples serve to demonstrate the fact that the learned dynamical model is (and should be) independent of the initial conditions of training trajectories. These solutions are respectively derived with $100\%$ and $10\%$ training data densities, both corrupted by $10\%$-level noise. The results present a close alignment between the predictions and ground truth of the trajectories.
To show the advantage of the proposed method, a comparison of relative errors in estimated parameters by Gaussian process learning and those by linear regression with finite-difference derivative approximation is depicted in Figure 3. The relative error $\epsilon_{1}$ was defined in (26). Both models are trained with 100%, 10%, 5%, and 1% data under 0%, 10% and 20% noise levels. Here, noisy data are smoothed by the Savitzky-Golay method [18]. Though the linear regression with finite-difference derivatives exhibits comparable accuracy to the proposed method when 100% noise-free data are used for training, it is consistently outperformed by the proposed method when noise is added or data are reduced.
Figure 2: Trajectory predictions for two new initial conditions issued by dynamical models respectively trained with $100\%$ and $10\%$ of data (2000 and 200 data points, respectively), both corrupted by $10\%$ noise.
Figure 3: Comparison of relative errors in the estimated parameters by the proposed Gaussian process learning (GPL) and those by linear regression with finite-difference derivative approximations (FD+LinReg). Both models are trained with 100%, 10%, 5% and 1% data (2000, 200, 100, and 20 points, respectively) under 0%, 10%, and 20% noise levels. In the latter case, noisy data are smoothed by the Savitzky-Golay method [18].
@@
Figure 4: Posterior distributions of ${\bf\it\theta}_{i}$ estimated under 0% and 20% noise levels with 10% and 50% data (200 and 1000 data points, respectively).
@@
Figure 6: Comparison of relative errors in the estimated parameters by the proposed Gaussian process learning (GPL) and those by SINDy. Both models are trained with 100%, 10%, 5%, and 1% data (2000, 200, 100, and 20 data points, respectively) under 0%, 10%, and 20% noise levels. In the noisy cases, the ensemble SINDy method with a bragging scheme is employed [15].
@@
The relative error in the posterior means and the relative deviation of the estimated parameters at different noise levels and amounts of training data, as previously defined in (26), are shown in Figure 5. It is evident that, overall, the values of both these error/uncertainty indicators are effectively decreased as the data amount is increased or the noise is reduced, i.e., an convergence is confirmed by these observations.
Besides, we present a comparison of relative errors between the proposed Gaussian process learning and SINDy, as shown in Figure 5. Both models are trained with 100%, 10%, 5%, and 1% data under 0%, 10% and 20% noise levels. In the noisy cases, ensemble SINDy with a bragging scheme is employed [15] to improve the performance of SINDy under uncertainties. We have observed that the proposed Gaussian-process-based method outperformed the SINDy methods in terms of estimation accuracy, particularly when dealing with noisy and sparse data.
@@
Results with scarce, noise-free data, including the trajectory reconstruction and prediction, as well as the approximation of $f(\cdot;{\bf\it\theta})$, are presented in Figure 7. The neural network approximates $f(\cdot;{\bf\it\theta})$ well within the range covered by training data, but fails in extrapolation when the prediction of $x$ goes beyond the training coverage. In fact, this observation is aligned with the limitation of neural network regression in generalization. In contrast to the previously discussed scenarios where the dictionary $\boldsymbol{g}_{i}(\boldsymbol{x})$ of candidate terms are considered to cover the expression of $f_{i}(\boldsymbol{x};\cdot)$, the neural network ansatz space here does not exactly include the actual underlying system representation but provides a reparametrization/approximation. This is commonly considered in the cases with unspecified nonlinear parametrization. Therefore, the poor predictive performance beyond training coverage is reasonable and expected. Nevertheless, such inadequacy in predictions can be effectively indicated by the uncertainty bands in our results. On the other hand, if the training data have covered the entire time domain of interest, the neural network reparametrization $f(\cdot;{\bf\it\theta})$ serves a surrogate model. As shown in Figure 8, only with 10 data points, the surrogate modeling is already decent and enables a good trajectory reconstruction.